# Learning Macroeconomic Policies based on Microfoundations: A Stackelberg Mean Field Game Approach

### v3

@@ 4 Stackelberg Mean Field Reinforcement Learning
To solve macroeconomic policymaking under Dynamic SMFG framework, we propose the Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm, which is a centralized training with decentralized execution (CTDE) algorithm. We introduce the leader’s policy $\pi^{l}(s^{l})$ and a central critic Q-function $Q^{l}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ for the leader agent, and the shared policy $\pi^{f}(s^{f},a^{l})$ and $Q^{f}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ for each follower agent. SMFRL algorithm comprises two key components: mean field approximation and leader-follower update (see Figure 2).
@@
Mean Field Approximation
Due to the dimensions of joint state $\mathbf{s}^{f}$ and action $\mathbf{a}^{f}$ scaling with the number of follower agents, the standard Q-function $Q^{l}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ becomes infeasible to learn when dealing with a large number of followers. Therefore, we factorize the standard Q-function using pairwise local interactions between the leader and each follower (Yang et al. 2018):
$Q^{l}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ $=\frac{1}{N}\sum_{i}^{N}\tilde{Q}^{l}(s^{l},a^{l},s^{fi},a^{fi})$ (2)
$\approx\mathbb{E}_{(s^{f},a^{f})\sim L}\tilde{Q}^{l}(s^{l},a^{l},s^{f},a^{f})$
Similarly, the $i$-th follower’s Q-function $Q^{f}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ can be factorized using local interactions between the leader, the $i$-th follower, and each of the remaining followers:
$Q^{f}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})\approx\mathbb{E}_{(s^{f},a^{f})\sim L}\tilde{Q}^{f}(s^{l},a^{l},s^{fi},a^{fi},s^{f},a^{f})$
where $(s^{fi},a^{fi})$ and $(s^{f},a^{f})$ are the state-action pair for the $i$-th follower and another follower agent, respectively. We use $\hat{L}(\pi^{f},a^{l})$, a finite set of state-action pairs sampled under the follower’s policy $\pi^{f}$ and the leader’s action $a^{l}$, to construct an empirical distribution that approximates the state-action distribution $L(s^{f},a^{f})$.
$\hat{L}(\pi^{f},a^{l})=\left\{(s^{fi},a^{fi})\right\}_{i=1}^{N},\text{ where }a^{fi}=\pi^{f}(s^{fi},a^{l}).$
Leader-follower Update
To enhance the convergence and stability, we propose the leader-follower update: first, by fixing the leader’s policy, we train the followers’ shared policy and Q-networks towards the best response; subsequently, based on the followers’ policies, we optimize the leader’s policy and Q net, alternating these steps until convergence. We measure the distance between the agents’ policies and their best responses by using exploitability.
Based on mean field approximation, we will train these networks $\pi_{\theta^{l}}$ and $\tilde{Q}_{\phi^{l}}$ for the leader agent, shared $\pi_{\theta^{f}}$ and $\tilde{Q}_{\phi^{f}}$ for the follower agents, with parameters $\theta^{l}$, $\phi^{l}$, $\theta^{f}$, and $\phi^{f}$. To ensure training stability, we introduce target networks with parameters $\theta^{l}_{-}$, $\phi^{l}_{-}$, $\theta^{f}_{-}$, and $\phi^{f}_{-}$. At any step $t\in\{0,\ldots,T\}$, the tuple $(s^{l}_{t},a^{l}_{t},\mathbf{s}^{f}_{t},\mathbf{a}^{f}_{t},s^{l}_{t+1},\mathbf{s}^{f}_{t+1},r^{l}_{t},\mathbf{r}^{f}_{t})$ is stored in the replay buffer $\mathcal{D}$ for training. When the number of followers is particularly large, a finite subset of state-action pairs $\{(s^{fi}_{t},a^{fi}_{t})\}_{i=1}^{N}$ can be selected for storage.
@@
Given leader’s policy $\pi_{\theta^{l}}$, the followers’ policy network $\pi_{\theta^{f}}$ is updated using a deterministic policy gradient approach (Silver et al. 2014). The followers’ policy gradient is estimated:
$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t}\sim\mathcal{D},(s^{f}_{t},a^{f}_{t})\sim\hat{L}_{\mathcal{D}}}$
$\Big{[}\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{fi}_{t},a^{l}_{t})\nabla_{a^{fi}_{-}}\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{fi}_{t},a^{fi}_{-},s^{f}_{t},a^{f}_{t})\Big{]}$
where $a^{fi}_{-}=\pi_{\theta^{f}}(s^{fi}_{t},a^{l}_{t})$, $\hat{L}_{\mathcal{D}}$ denotes state-action pairs sampled from replay buffer $\mathcal{D}$. The action-value function $\tilde{Q}_{\phi^{f}}$ is updated by minimizing the mean squared error loss:
$\mathcal{L}(\phi^{f})=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t},a^{l}_{t},\mathbf{s}_{t+1}\sim\mathcal{D}\\ (s^{f}_{t},a^{f}_{t})\sim\hat{L}_{\mathcal{D}}\end{subarray}}\left[\left(y^{fi}_{t}-\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{fi}_{t},a^{fi}_{t},s^{f}_{t},a^{f}_{t})\right)^{2}\right]$
$y^{fi}_{t}=r^{fi}_{t}+$
$\gamma\mathbb{E}_{(s^{f}_{t+1},a^{f}_{t+1})\sim\hat{L}(\pi_{\theta^{f}_{-}},a^{l}_{t+1})}\tilde{Q}_{\phi^{f}_{-}}(s^{l}_{t+1},a^{l}_{t+1},s^{fi}_{t+1},a^{fi}_{t+1},s^{f}_{t+1},a^{f}_{t+1})$
where $a^{l}_{t+1}=\pi_{\theta^{l}}(s^{l}_{t+1})$, $a^{fi}_{t+1}=\pi_{\theta^{f}_{-}}(s^{fi}_{t+1},a^{l}_{t+1})$. The gradient of the loss function $\mathcal{L}(\phi^{f})$ is derived as:
$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t},a^{l}_{t},\mathbf{s}_{t+1}\sim\mathcal{D}\\ (s^{f}_{t},a^{f}_{t})\sim\hat{L}_{\mathcal{D}}\end{subarray}}$ $\Big{[}\left(y^{fi}_{t}-\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{fi}_{t},a^{fi}_{t},s^{f}_{t},a^{f}_{t})\right)$
$\nabla_{\phi^{f}}\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{fi}_{t},a^{fi}_{t},s^{f}_{t},a^{f}_{t})\Big{]}.$
@@
$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t}\sim\mathcal{D},(s^{f},a^{f})\sim\hat{L}(\pi_{\theta^{f}},a^{l}_{-})}$
$\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}\tilde{Q}_{\phi^{l}}(s^{l},a^{l}_{-},s^{f},a^{f})\right],\text{where }{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}.$
@@
$\mathcal{L}(\phi^{l})=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t},a^{l}_{t},\mathbf{s}_{t+1}\sim\mathcal{D}\\ (s^{f}_{t},a^{f}_{t})\sim\hat{L}_{\mathcal{D}}\end{subarray}}\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},s^{f}_{t},a^{f}_{t})\right)^{2}\right]$
@@
$y^{l}_{t}=r_{t}^{l}+\gamma\mathbb{E}_{(s^{f}_{t+1},a^{f}_{t+1})\sim\hat{L}(\pi_{\theta^{f}},a^{l}_{t+1})}\tilde{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},s^{f}_{t+1},a^{f}_{t+1})$
where $a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})$, and $\gamma$ is the discount factor. Differentiating the loss function $\mathcal{L}(\phi^{l})$ yields the gradient utilized for training:
$\nabla_{\phi^{l}}$ $\mathcal{L}(\phi^{l})=\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},\mathbf{s}_{t+1}\sim\mathcal{D},(s^{f}_{t},a^{f}_{t})\sim\hat{L}_{\mathcal{D}}}$
$\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},s^{f}_{t},a^{f}_{t})\right)\nabla_{\phi^{l}}\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},s^{f}_{t},a^{f}_{t})\right].$
The pseudocode for the SMFRL algorithm 1 is presented in Appendix A.
@@ 5 Experiment
In this section, we conduct experiments of the SMFG method, which is modeled by dynamic SMFGs and solved by the SMFRL algorithm, within the TaxAI (Mi et al. 2023) environment to answer the following questions:
1. For macroeconomic policy-making issues, how does our SMFG method compare to real-world policies, classic economic methods, and existing AI-based policy?
2. Are both the leader-follower update and mean-field approximation modules of the SMFG method essential?
3. Given the likely real-world scenario where some households do not adopt the SMFG policy, does the SMFG approach remain effective?
TaxAI Environment
TaxAI is a simulation environment based on real data designed to study optimal taxation policies. It simulates the dynamic interactions between the government and large-scale households, to promote economic growth (i.e., GDP) as the objective.
@@
Tax Policies Per Capita GDP Social Welfare Income Gini Wealth Gini Years
$100$ $1000$ $100$ $1000$ $100$ $1000$ $100$ $1000$ $100$ $1000$
Free Market 1.37e+05 1.41e+05 32.97 334.79 0.89 0.90 0.92 0.93 1.10 1.00
Saez Tax 2.34e+12 6.35e+11 73.82 498.88 0.21 0.68 0.38 0.73 300.00 100.58
US Federal Tax 4.88e+11 1.41e+05 94.19 351.17 0.40 0.89 0.40 0.93 289.55 1.00
AI Economist-BC 4.24e+12 N/A 97.24 N/A 0.54 N/A 0.52 N/A 299.55 N/A
AI Economist 1.26e+05 N/A 72.81 N/A 0.88 N/A 0.91 N/A 1.00 N/A
SMFG 9.59e+12 1.10e+13 96.87 968.94 0.52 0.54 0.51 0.53 300.00 300.00
Table 1: Assessment results of 6 macroeconomic policies for different numbers of households N=100 and N=1000.
Figure 3: Dynamic response curves of 3 macroeconomic policies to economic shocks at step 100. The SMFG policy (dark blue line) exhibits the least fluctuation and the fastest recovery in key indicators, indicating superior dynamic response capabilities.
Figure 4: Ablation studies of SMFG method based on macroeconomic indicators for households’ numbers N=100 and N=1000.
Comparison of Macroeconomic Policies
This section compares the tax policy derived from SMFG with existing taxation policies in two aspects: (1) Macroeconomic Policy Evaluation: We assess the performance of different policies using key macroeconomic indicators. (2) Dynamic Response: At step 100, we simulate an economic shock caused by a financial crisis to evaluate the dynamic response capabilities of the different tax policies.
Baselines and Metrics
We select four different types of tax policy baselines, including the Free Market (Backhouse 2005) policy, which entails no government intervention; the Saez Tax (Saez 2001), a well-known economic approach often advocated for specific tax reforms in practice; the U.S. Federal Tax, which is the actual policy derived from OECD data https://data.oecd.org/; and the AI Economist (Zheng et al. ), a popular AI-based policy utilizing a two-level Proximal Policy Optimization (PPO) framework. Additionally, the AI Economist-BC method employs behavior cloning on real-world data for pretraining. In the TaxAI environment, we evaluate the performance of these policies using key macroeconomic indicators derived from real-world data. Per Capita GDP reflects the level of economic development, while Income Gini and Wealth Gini measure inequality in household income and wealth, respectively—a lower Gini index indicates greater social equality. The Years metric represents the sustainable duration of an economy, with a maximum cap of 300 years. Average Wealth, Income, and Consumption are crucial assessment metrics related to financial crises.
Macroeconomic Policy Evaluation
In the context of optimizing economic growth, our SMFG method achieves the highest per capita GDP, social welfare, and the longest duration of outcomes across both $N=100$ and $N=1000$ scenarios. Table 1 shows the average values of macroeconomic indicators across 10 test runs for different algorithms under varying household numbers. Although the Gini index for SMFG is higher than that of the Saez Tax when $N=100$, this is acceptable when optimizing GDP, as SMFG’s per capita GDP is nearly 4 times that of the Saez Tax. When $N=1000$, SMFG outperforms all other metrics, indicating that the advantages of our SMFG method increase with the number of households. Figure 7 in the Appendix provides the corresponding training curves.
Dynamic Response
We simulate an economic shock similar to a financial crisis in the TaxAI environment: at step 100, all households lose 50% of their wealth. Due to the short duration (Years) under the Free Market and AI economist strategies, we focus on comparing the Saez Tax, US Federal Tax, and SMFG policy.
From Figure 3, we find that the SMFG policy (dark blue line) demonstrates the fastest dynamic response across all indicators: Social Welfare remains largely unaffected, while Average Income and Average Consumption recover to their original levels within about 5 steps. Wealth Gini and Average Wealth return to their previous levels within about 15 steps. This level of recovery is unmatched by any other tax policy.
Figure 5: The left subfigure illustrates a scenario with heterogeneous households adopting various policies. The right subfigure presents test results for households’ decisions (consumption, labor), microeconomic indicators (utility, wealth, income), and macroeconomic indicators (GDP) when there are different proportions of SMFG households ($N=100$).
Ablation Studies of SMFG
We conduct ablation studies to analyze the impact of the leader-follower update and mean field approximation modules in the SMFG method on macroeconomic indicators, as well as on convergence and social welfare.

### v4

@@ 4 Stackelberg Mean Field Reinforcement Learning
Within the DSMFG framework, we propose the Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm (Figure 2), which adopts a centralized training with decentralized execution (CTDE) paradigm [36]. In standard CTDE settings, the leader agent is equipped with a policy $\pi^{l}(s^{l})$ and a centralized critic $Q^{l}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$, while each follower is equipped with a policy $\pi^{f}(s^{f},a^{l})$ and a corresponding Q-function $Q^{f}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$. However, the joint state $\mathbf{s}^{f}$ and action $\mathbf{a}^{f}$ of all followers scale linearly with the population size, rendering direct learning of $Q^{l}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ computationally infeasible in large-scale environments. This directly reflects the computational complexity challenge induced by large-scale agent interactions, as discussed in Section 3.
4.1 Stackelberg Mean-Field Q and Policy
Stackelberg Mean-Field Q
Within the DSMFG framework, we abstract the interactions between the leader and all followers into an interaction between the leader and the population mean field, thereby the leader’s centralized Q-function can be reformulated as:
$Q^{l}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})\approx\tilde{Q}^{l}(s^{l},a^{l},L).$
Similarly, based on the mean field approximation, we simplify the interaction of a follower with the leader and other followers into an interaction with the leader and the population mean field. Accordingly, the original Q-function $Q^{f}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ can be approximated as:
$Q^{f}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ $=Q^{f}(s^{l},a^{l},s^{fi},a^{fi},\mathbf{s}^{f,-i},\mathbf{a}^{f,-i})$
$\approx\tilde{Q}^{f}(s^{l},a^{l},s^{fi},a^{fi},L),$
where $(s^{fi},a^{fi})$ denotes the individual state-action pair of the $i$-th follower, and $(\mathbf{s}^{f,-i},\mathbf{a}^{f,-i})$ represent the rest of the population. Given that each individual has a negligible impact on the collective, the population mean field $L$ can be used to approximate the rest of the population, following the theory of mean field games [33].
In experiments, the mean field $L$ is constructed from the empirical distribution over the followers’ state-action pairs. At timestep $t$,
$L_{t}=P(s^{f},a^{f}),\quad(s^{f},a^{f})\sim\{(s^{f_{i}}_{t},a^{f_{i}}_{t})\}_{i=1}^{N},a^{f_{i}}_{t}\sim\pi^{f}(\cdot\mid s^{f_{i}}_{t},a^{l}_{t}),$
where $N$ denotes the number of followers.
In prior works on mean-field methods, a widely used simplification is to approximate $L_{t}$ using population averages:
$L_{t}\approx(\bar{s}^{f},\bar{a}^{f}),\quad\bar{s}^{f}=\frac{1}{N}\sum_{i=1}^{N}s^{f_{i}},\quad\bar{a}^{f}=\frac{1}{N}\sum_{i=1}^{N}a^{f_{i}}.$ (2)
Alternative representations include neighborhood-based action averages [59], empirical distributions [13], and graph-based weighted mean fields [24]. In our experiments, we find that using the average-based mean field achieves strong empirical performance while maintaining tractable computational cost.
In reinforcement learning, the local Q-functions $\tilde{Q}^{l}$ and $\tilde{Q}^{f}$ are defined as the expected cumulative rewards under discount factor $\gamma\in[0,1]$, starting from given states and actions:
$\tilde{Q}^{l}(s^{l},a^{l},s^{f},a^{f})=\mathbb{E}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}^{l}\right],$ (3)
$\tilde{Q}^{f}(s^{l},a^{l},s^{fi},a^{fi},s^{f},a^{f})=\mathbb{E}\left[\sum_{t=0}^{T}\gamma^{t}r^{fi}_{t}\right].$
Here, $r_{t}^{l}$ and $r_{t}^{fi}$ are the rewards for the leader and $i$-th follower, respectively. The discount factor $\gamma$ determines the planning horizon: $\gamma=0$ models myopic agents focusing on immediate rewards, while $\gamma>0$ models non-myopic agents with long-term effects, increasing computational complexity.
Followers’ policy
Under the standard CTDE paradigm, the leader’s policy takes the individual state as input, while the follower’s policy conditions on both the individual’s state and the leader’s action. These input dimensions are tractable. However, training a separate policy for each follower is computationally infeasible.
To address this, we adopt a shared policy $\pi^{f}(s^{f},a^{l})$ for all followers, as required by the mean-field setting. While this introduces inherent homogeneity—a known limitation of mean-field methods—we preserve individual heterogeneity by encoding personalized information in the state input.
In reinforcement learning, a policy maps states to actions. Through training, the model learns how state features influence decisions. For instance, individual states include attributes such as age, education, and wealth, while actions cover economic choices like investment. This enables the shared policy to generate personalized behaviors, e.g., individuals of different ages and wealth levels exhibit distinct investment patterns.
@@ 4.1 Stackelberg Mean-Field Q and Policy
4.2 Leader-follower Update
To enhance the convergence and stability, we propose the leader-follower update (shown in Figure 2): first, by fixing the leader’s policy, we train the followers’ shared policy and Q-networks towards the best response; subsequently, based on the followers’ policies, we optimize the leader’s policy and Q net, alternating these steps until convergence. We measure the distance between the agents’ policies and their best responses by using exploitability.
Based on mean field approximation, we will train these networks $\pi_{\theta^{l}}$ and $\tilde{Q}_{\phi^{l}}$ for the leader agent, shared $\pi_{\theta^{f}}$ and $\tilde{Q}_{\phi^{f}}$ for the follower agents, with parameters $\theta^{l}$, $\phi^{l}$, $\theta^{f}$, and $\phi^{f}$. To ensure training stability, we introduce target networks with parameters $\theta^{l}_{-}$, $\phi^{l}_{-}$, $\theta^{f}_{-}$, and $\phi^{f}_{-}$. At any step $t\in\{0,\ldots,T\}$, the tuple $(s^{l}_{t},a^{l}_{t},\mathbf{s}^{f}_{t},\mathbf{a}^{f}_{t},s^{l}_{t+1},\mathbf{s}^{f}_{t+1},r^{l}_{t},\mathbf{r}^{f}_{t})$ is stored in the replay buffer $\mathcal{D}$ for training.
@@ 4.2 Leader-follower Update
Given leader’s policy $\pi_{\theta^{l}}$, the followers’ policy network $\pi_{\theta^{f}}$ is updated using a deterministic policy gradient [54]. The followers’ policy gradient is estimated:
$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},L_{t}\sim\mathcal{D}}\Big{[}\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})\nabla_{a^{f}_{-}}\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{f}_{t},a^{f}_{-},L_{t})\Big{]}$
where $a^{f}_{-}=\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})$, $L_{t}$ is computed by state-action pairs sampled from replay buffer $\mathcal{D}$ by Eq. (2). The action-value function $\tilde{Q}_{\phi^{f}}$ is updated by minimizing the mean squared error loss:
$\mathcal{L}(\phi^{f})=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t},a^{l}_{t},L_{t},\mathbf{s}_{t+1}\sim\mathcal{D}\end{subarray}}\left[\left(y^{f}_{t}-\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{f}_{t},a^{f}_{-},L_{t})\right)^{2}\right]$
$y^{f}_{t}=r^{f}_{t}+\gamma\tilde{Q}_{\phi^{f}_{-}}(s^{l}_{t+1},a^{l}_{t+1},s^{f}_{t+1},a^{f}_{t+1},L_{t+1})$
where $a^{l}_{t+1}=\pi_{\theta^{l}}(s^{l}_{t+1})$, $a^{f}_{t+1}=\pi_{\theta^{f}_{-}}(s^{f}_{t+1},a^{l}_{t+1})$. The gradient of the loss function $\mathcal{L}(\phi^{f})$ is derived as:
$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t},a^{l}_{t},L_{t},\mathbf{s}_{t+1}\sim\mathcal{D}\end{subarray}}$ $\Big{[}\left(y^{f}_{t}-\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{f}_{t},a^{f}_{t},L_{t})\right)$
$\nabla_{\phi^{f}}\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{f}_{t},a^{f}_{t},L_{t})\Big{]}.$
@@
$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t},L_{t}\sim\mathcal{D}}\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}\tilde{Q}_{\phi^{l}}(s_{t}^{l},a^{l}_{-},L_{t})\right]|_{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}.$
@@
$\mathcal{L}(\phi^{l})=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t},a^{l}_{t},L_{t},\mathbf{s}_{t+1}\sim\mathcal{D}\end{subarray}}\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L_{t})\right)^{2}\right]$
@@
$y^{l}_{t}=r_{t}^{l}+\gamma\tilde{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},L_{t+1})|_{a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})}$
where $\gamma$ is the discount factor. Differentiating the loss function $\mathcal{L}(\phi^{l})$ yields the gradient utilized for training:
$\nabla_{\phi^{l}}\mathcal{L}(\phi^{l})=\mathbb{E}\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L_{t})\right)\nabla_{\phi^{l}}\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L_{t})\right].$
where the expectation $\mathbb{E}$ is taken over $(\mathbf{s}_{t},a^{l}_{t},L_{t},\mathbf{s}_{t+1})\sim\mathcal{D}$. The pseudocode for the SMFRL algorithm 1 in Appendix A.
@@ 5 Experiment
Table 1: Performance of multiple policies on key macroeconomic indicators for $N=100$ and $N=1000$ households. The best values are highlighted in bold, and the second-best values are underlined.
@@
Category Subcategory Policies Per Capita GDP $\uparrow$ Social Welfare $\uparrow$ Wealth Gini $\downarrow$ Years $\uparrow$
$100$ $1000$ $100$ $1000$ $100$ $1000$ $100$ $1000$
Static Non-intervention Free Market 1.37e+05 1.41e+05 32.97 334.79 0.92 0.93 1.10 1.00
Policies Real-data US Federal Tax 4.88e+11 1.41e+05 94.19 351.17 0.40 0.93 289.55 1.00
Dynamic Rule-based Saez Tax 2.34e+12 6.35e+11 73.82 498.88 0.38 0.73 300.00 100.58
Policies Independent-based AI Economist 1.26e+05 N/A 72.81 N/A 0.91 N/A 1.00 N/A
Game-based DSMFG (ours) 9.59e+12 1.10e+13 96.87 968.94 0.51 0.53 300.00 300.00
DSMFG w/o S 8.66e+07 1.31e+05 82.02 834.93 0.83 0.92 75.75 1.50
DSMFG w/o MF 1.23e+05 1.48e+05 48.17 499.01 0.93 0.93 1.00 1.02
DSMFG w/o MF & S 1.21e+05 1.33e+05 83.09 874.53 0.92 0.92 1.00 1.00
To validate the effectiveness of our DSMFG framework and SMFRL algorithm for macroeconomic policymaking, our experiments are designed to answer the following key questions:
  1. 1.
    Effectiveness of dynamic modeling. Does the DSMFG framework outperform static macroeconomic policies in optimizing critical economic indicators? (§ 5.2)
  2. 2.
    Necessity of Stackelberg structure and mean field approximation. Are the Stackelberg structure and mean field approximation essential for the scalability and performance of DSMFG? (§ 5.3)
  3. 3.
    Impact of mean field homogeneity. How does the homogeneity assumption inherent in mean field approximations affect decision personalization and policy robustness? (§ 5.4)
In the appendix, we include details on computational resources and efficiency (E.1), full training curves and result tables (E.2,E.3), and discussions on the efficiency-equity of different policies (E.4).
5.1 Experimental Setting
Environment
We conduct experiments in TaxAI [39], a simulation platform for optimal tax policy. TaxAI enables dynamic interactions between governments and large-scale households using real-world datasets. Details are in Appendix F.1.
Evaluation Metrics
Per Capita GDP reflects the level of economic development, while Income Gini and Wealth Gini measure inequality in household income and wealth, respectively—a lower Gini index indicates greater social equality. The Years metric represents the sustainable duration of an economy, with a maximum cap of 300 years. Average Wealth, Income, and Consumption are crucial assessment metrics related to financial crises.